28th September (Monday)

  • Data Processing:
    • Data processing of pp and pHe data at T0/1 sites, monte carlo mostly at T2, user analysis at T0/1/2D sites

  • T0
    • Investigation of slow worker ongoing is followed up internally the ticket is closed (GGUS:116023)
    • Problem accessing EOS caused by Namespace service stuck (GGUS:116445), issue solved
    • Failed transfers to CERN-RAW, re-surfaced on Sunday (GGUS:116321)
    • Observed that pilots seem to get stuck in the CREAM-CEs of LSF and not submitted to batch (GGUS:116473)

24th September (Thursday)

  • Data Processing:
    • Data processing of the proton-proton and proton-helium data has been started.
    • Data processing, MC and user jobs

  • T0
    • Investigation of slow worker ongoing is followed up internally the ticket is closed (GGUS:116023)
    • Permission denied on eos for some users (in progress) (GGUS:116243)
    • Failed transfers to CERN-RAW (GGUS:116321)

17th September (Thursday)

  • Data Processing:
    • Validation of data reconstruction, MC and user jobs

14th September (Monday)

  • Data Processing:
    • Dominated by MC and user jobs

  • T0

11th September (Friday)

(Apologies for not being able to deliver this report in person/over Vidyo!)

  • Data Processing:
    • Dominated by MC and user jobs

  • T0
    • Investigation of slow worker ongoing (GGUS:116023)
    • dbod 9 Sep downtime - DB transfer complete and all services restarted without issue

  • T1
    • CNAF: Network issues causing slow downloads/uploads (GGUS:116023)

7th September (Monday)

  • Data Processing:
    • Dominated by MC and user jobs

  • T0
    • Investigation of slow worker ongoing (GGUS:116023)
    • dbod 9 Sep downtime - Plans in place, waiting for final confirmation about cluster migration beforehand.

  • T1
    • CNAF Outage: All seems OK now
    • SARA: DNS intervention caused some failed transfers over the weekend. All OK now.

2nd September (Tuesday)

  • Data Processing:
    • Dominated by MC and user jobs

  • T0
    • identified another slow worker node (GGUS:116023)
    • dbod 9 Sep downtime - asking for move to cluster installation before
    • lbvobox downtime 11 Sep - is it possible to do live migration of a VM?

  • T1
    • RAL: ARC-CE submission working again
    • CNAF OUTAGE downtime declared Friday evening for LHCb storage

-- JoelClosier - 2016-01-07

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2016-01-07 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback